### v1

# Zaragoza, Spain Feature Splatting

Abstract
3D Gaussian Splatting has emerged as a very promising scene representation, achieving state-of-the-art quality in novel view synthesis significantly faster than competing alternatives. However, its use of spherical harmonics to represent scene colors limits the expressivity of 3D Gaussians and, as a consequence, the capability of the representation to generalize as we move away from the training views. In this paper, we propose to encode the color information of 3D Gaussians into per-Gaussian feature vectors, which we denote as Feature Splatting (FeatSplat). To synthesize a novel view, Gaussians are first "splatted" into the image plane, then the corresponding feature vectors are alpha-blended, and finally the blended vector is decoded by a small MLP to render the RGB pixel values. To further inform the model, we concatenate a camera embedding to the blended feature vector, to condition the decoding also on the viewpoint information. Our experiments show that these novel model for encoding the radiance considerably improves novel view synthesis for low overlap views that are distant from the training views. Finally, we also show the capacity and convenience of our feature vector representation, demonstrating its capability not only to generate RGB values for novel views, but also their per-pixel semantic labels. We will release the code upon acceptance.
Keywords: Gaussian Splatting, Novel View Synthesis, Feature Splatting
1 Introduction
Finding appropriate 3D scene representations is a key challenge in order to enable robotics, VR and AR applications. In general, representations should not only encode a scene’s geometry. Radiance properties are very relevant for novel view synthesis and high-level aspects such as semantics or affordances are essential to define robotics tasks. In the last years, Neural Radiance Fields (NeRFs) [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] have been the dominant approach in the research literature to learn implicit scene representations from images. However, NeRFs are computationally demanding at training and rendering and scale badly with the size of the scene, which has limited their use in practical applications.
3D Gaussian Splatting (3DGS) [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] appeared very recently as a promising alternative to NeRFs. 3DGS is able to achieve similar quality for novel view synthesis significantly faster, although with a much higher memory footprint. Representing a scene with 3D Gaussians, however, does not come without limitations. As the most relevant to our work, a explicit scene representation composed of 3D Gaussians and spherical harmonics lacks the capacity to model complex image textures. This limitation is typically alleviated by an increase of the number of Gaussians and a reduction of their size, so that each tiny Gaussian models a smaller and simpler part of the texture. Such strategy, in addition to being intensive in memory, leads to overfitting to the training views and to a poor generalization.
In this paper we propose a novel approach, that we denote as Feature Splatting (FeatSplat), that replaces spherical harmonics by a feature vector representation for each 3D Gaussian. Using such representations, we show in ScanNet++ [Yeshwanth et al.(2023)Yeshwanth, Liu, Nießner, and Dai] a significant improvement in novel view synthesis with low overlap. This comes as a result of the higher capacity of our feature vector representation compared to spherical harmonics. Figure 1 illustrates our method. We splat the 3D Gaussians and alpha-blend the feature vectors. We then concatenate to the splatted and blended feature vector dimensions to camera intrinsics and extrinsics channels, and finally decode them into colors with a MLP. As an additional use case illustrating the flexibility of our representation, we also show its capability to render accurate semantic labels.
2 Related Work
In the last years, fully implicit scenes representations have been typically modelled as a function that maps scene coordinates to the desired scene attributes (e.g., color), and a neural network is trained on the scene data so that it learns such function. One of the most successful examples are NeRFs [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng]. A NeRF consists on a Multi Layer Perceptron (MLP) that maps a 3D point and a viewing direction to a RGB color and an occupancy or density value, coupled with a differentiable rendering equation to aggregate values of 3D points along a ray casted through each pixel to compute the final RGB color. Given a set of posed images, the MLP is optimized to perform novel view synthesis afterwards. To compensate for the initial high training and rendering times, several approaches have been developed. Some proposed to use depth information to reduce sampling in 3D [Deng et al.(2022)Deng, Liu, Zhu, and Ramanan, Wei et al.(2021)Wei, Liu, Rao, Zhao, Lu, and Zhou, Neff et al.(2021)Neff, Stadlbauer, Parger, Kurz, Mueller, Chaitanya, Kaplanyan, and Steinberger, Roessle et al.(2022)Roessle, Barron, Mildenhall, Srinivasan, and Nießner]; others combined the MLP with explicit 3D representations like Octrees [Yu et al.(2021)Yu, Li, Tancik, Li, Ng, and Kanazawa, Fridovich-Keil et al.(2022)Fridovich-Keil, Yu, Tancik, Chen, Recht, and Kanazawa], Voxels [Liu et al.(2020)Liu, Gu, Zaw Lin, Chua, and Theobalt], 3D point clouds [Xu et al.(2022)Xu, Xu, Philip, Bi, Shu, Sunkavalli, and Neumann] and 3D hierarchical mesh grids [Müller et al.(2022)Müller, Evans, Schied, and Keller] and encoded into it learned features that where then quickly decoded using a smaller MLP.
Recently, 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] has emerged as a substantially faster yet still high-quality alternative. It relies on the explicit modelling of the 3D geometry by a set of 3D Gaussians, each of them with a set of spherical harmonics (SHs) [Yu et al.(2021)Yu, Li, Tancik, Li, Ng, and Kanazawa] to represent their color. Rather than using ray tracing like NeRFs, Gaussians are rasterized by being projected into the image plane, sorted by the Radix algorithm [Merrill and Grimshaw(2010)], and combined using alpha-blending to give the RGB color for every pixel. The rendered image is then used to compute a photometric loss, and the Gaussians’ parameters are optimized using backpropagation. By avoiding the MLP training and relying in an efficient sorting algorithm and CUDA implementation of the rasterizer, 3DGS is the state of the art for the combined rendering performance, training time, and inference rendering speed. Nevertheless, 3DGS is also characterized by 1) a high memory usage; 2) poor surfaces’ reconstructions; and 3) the limited capacity of SHs, which increases the number of Gaussians needed for high quality results, worsening the previous two problems. Several approaches augment Gaussians adding latent vectors which are latter decoded. LangSplat [Qin et al.(2023)Qin, Li, Zhou, Wang, and Pfister] combines CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] features with SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] segmentation masks to generate a language field enabling open-vocabulary 2D semantic segmentation. FMGS [Zuo et al.(2024)Zuo, Samangouei, Zhou, Di, and Li] uses a hash grid based on iNGP [Müller et al.(2022)Müller, Evans, Schied, and Keller] to encode CLIP and DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, Howes, Huang, Xu, Sharma, Li, Galuba, Rabbat, Assran, Ballas, Synnaeve, Misra, Jegou, Mairal, Labatut, Joulin, and Bojanowski] features, to also generate an intrinsic representation capable of open-vocabulary 2D segmentation. Nevertheless, none of them neither address the SHs capacity nor show improvement in novel view synthesis, as we do. Based also on iNGP, Compact-3DGS [Lee et al.(2023)Lee, Rho, Sun, Ko, and Park] adds a hash grid that encodes a latent vector for each Gaussian and uses a small MLP conditioned by the viewing direction to recover the corresponding color for each point-of-view. Similarly, our FeatSplat relies on using a latent vector to encode each Gaussian’s color. But, differently from it, we directly attach a feature vector to each Gaussian, and, during the rasterization, rather than first decoding the colors and then alpha-blending, we first alpha-blend the feature vectors of corresponding Gaussians, and then decode the final vector to RGB, which allows us to condition on the intrinsic and extrinsic camera parameters. In our experiments, we show that FeatSplat, due to these differences, outperforms Compact-3DGS at novel view synthesis.
Camera
  Projection
Refer to captionAdaptiveDensity ControlRefer to caption
Differentiable
   Rasterizer
SfM initialized
3D Gaussians
$G:\{\mathbf{x},R,S,\sigma,\mathbf{f}\}$
Operation flow
Gradient flow
MLP
Feature
 Image
Figure 1: Overview of our approach. Left: We augment the 3D Gaussian Splatting [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] representation with learned feature vectors $\mathbf{f}$ to encode colors, removing the spherical harmonics. Center: To render a point of view, 3D Gaussians are projected to the image plane, where the differentiable rasterizer alpha-blends corresponding feature vectors. Resulting vectors are concatenated with a camera embedding, and a tiny MLP renders the final RGB values, with potentially also higher level information such as semantic labels. Right: Illustrative result of novel view synthesis result with low overlap, with RGB values and semantic labels.
3 Preliminaries: 3D Gaussian Splatting
3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] is an explicit scene representation that relies on a set of 3D Gaussians to encode a scene, which is optimized by differentiable rasterization from known viewpoints.
Representation. The geometry of a 3D Gaussian $G$ is parameterized by its 3D center position $\mathbf{x}_{G}\in\mathbb{R}^{3}$ and a 3D covariance $\Sigma_{G}=R_{G}S_{G}S_{G}^{\top}R_{G}^{\top}\succeq 0$ that can be decomposed into a rotation matrix $R_{G}\in SO(3)$ and a scaling matrix $S_{G}\in\text{diag}(\mathbb{R}^{3}_{\geq 0})$. Each Gaussian has a density value ${\sigma}_{G}\in\mathbb{R}$; and a set of spherical harmonics (SHs) coeficients $\mathbf{k}_{G}\in\mathbb{R}^{h}$ that parameterize the function $\mathcal{S}_{\mathbf{k}_{G}}\left(\mathbf{d}_{G}\right)$ to encode RGB values $\mathbf{c}_{G}\in\mathbb{R}^{3}$ depending on the viewing direction $\mathbf{d}_{G}\in S^{2}$ relative to each Gaussian.
Differentiable rendering. To render an RGB image $\hat{{I}}\in\mathbb{R}^{w\times h\times 3}$ of the scene, first SHs are converted to RGB, $\mathbf{c}_{G}=\mathcal{S}_{\mathbf{k}_{G}}\left(\mathbf{d}_{G}\right)$. Then, Gaussians are projected, or splatted, into 2D Gaussians in the image plane [Zwicker et al.(2001)Zwicker, Pfister, Van Baar, and Gross] and sorted using the Radix algorithm. Finally, the color $\textbf{c}_{p}$ for each pixel $\mathbf{p}\in\Omega$, $\Omega$ standing for the image domain, comes from the alpha blending of the $\mathcal{N}\subset G$ 2D Gaussians falling into it
$\textbf{c}_{p}=\sum_{i\in\mathcal{N}}{T_{i}\alpha_{i}\mathbf{c}_{i}}\text{, with }T_{i}=\prod_{j=i}^{i-1}\left(1-\alpha_{i}\right),$ (1)
where $\alpha_{i}\in\mathbb{R}$ is obtained by multiplying the learned occupancy value $\sigma_{i}$ by the evaluation of the projected 2D Gaussian [Yifan et al.(2019)Yifan, Serena, Wu, Öztireli, and Sorkine-Hornung, Kopanas et al.(2021)Kopanas, Philip, Leimkühler, and Drettakis, Kopanas et al.(2022)Kopanas, Leimkühler, Rainer, Jambon, and Drettakis] for each pixel $\mathbf{p}$.
Optimization. The representation is fitted as follows. The 3D Gaussians are initialized either randomly or from a sparse Structure from Motion point cloud. At each optimization step, the Gaussians are used to render a viewpoint from the training set and compute the loss
$\mathcal{L}=\left(1-\lambda_{SSIM}\right)\mathcal{L}_{1}\left(\hat{{I}},{I}\right)+\lambda_{SSIM}\mathcal{L}_{D-SSIM}\left(\hat{{I}},{I}\right),$ (2)
between the rendered image $\hat{{I}}$ and the ground truth $I$. $\mathcal{L}_{1}(\cdot,\cdot)$ stands for the L1-norm of the difference of its two arguments, $\mathcal{L}_{D-SSIM}(\cdot,\cdot)$ is the Structure Similarity Index Measure (SSIM) [Wang et al.(2004)Wang, Bovik, Sheikh, and Simoncelli] between its two arguments, and $\lambda_{SSIM}$ weights both terms. Adam [Kingma and Ba(2015)] is used to update the Gaussians’ parameters by backpropagating the loss, and Gaussians are pruned or cloned adaptively according to certain policies.
Despite SHs encoding direction-dependent RGB values, their variability is limited, and are only able to encode one color per viewing direction. To represent complex textures in natural scenes, 3DGS over-densifies scenes by optimizing Gaussians at different depths, visible only from certain viewpoints. In a manner, this strategy abuses the geometric primitive to overfit the photometry, achieving high-quality rendering for views close to the training distribution, but causing artifacts observable from viewpoints far from the training views.
4 Feature Splatting (FeatSplat)
Our FeatSplat relies on learned feature vectors instead of SHs to better encode information and increase the capacity and flexibility of 3D Gaussians. Each 3D Gaussian is initialized with a feature vector $\mathbf{f}_{G}$ sampled from a normal distribution $\mathcal{N}\left(0,1\right)$. When rendering an image $\hat{I}$, the $\mathcal{N}$ $\mathbf{f}_{G}$ vectors splatted in a pixel $\mathbf{p}$ are alpha-blended
$\textbf{f}_{p}=\sum_{i\in\mathcal{N}}{T_{i}\alpha_{i}\mathbf{f}_{i}},$ (3)
similarly to what was done with colors in Eq. (1). After that, a small MLP transforms $\textbf{f}_{p}$, into the corresponding RGB color $\textbf{c}_{p}$ of the target image. The per-pixel feature vectors $\textbf{f}_{p}$ are concatenated with camera parameter embeddings, in order to condition the MLP not only on the information of the 3D Gaussians, but also on the viewpoint. The camera extrinsics are represented by the 3D camera position of the target image in global coordinates, and the intrinsics by an embedding of the relative position of the rendered pixel in the target image plane. The final rendered color is then
$\textbf{c}_{p}=\text{MLP}\left(\textbf{f}_{p}\oplus\mathbf{x}_{cam}\oplus\textbf{e}_{p}\right),$ (4)
where $\oplus$ is a concatenation operation, $\mathbf{x}_{cam}\in\mathbb{R}^{3}$ is the 3D camera position, and $\mathbf{e}_{p}\in[-1,1]\times[-1,1]$ the two-dimensional pixel embedding. We also considered adding camera orientation information, but our ablations (App. A.2) did not show a benefit.
Semantic segmentation. Learning per-Gaussian feature vectors $\textbf{f}_{p}$ increases the capability of a 3DGS representation. In this paper we illustrate how FeatSplat can be adapted for 2D semantic segmentation. All is needed is changing the output dimension of the MLP from $3$ channels to $3+\#\text{classes}$ and training with an additional semantic cross-entropy loss $\mathcal{L}_{ce}$.
$\mathcal{L}=\left(1-\lambda\right)\mathcal{L}_{1}\left(\hat{\text{I}},\text{I}\right)+\lambda\mathcal{L}_{D-SSIM}\left(\hat{\text{I}},\text{I}\right)+\lambda_{sem}\mathcal{L}_{ce}\left(\hat{S},S\right),$ (5)
where $\lambda_{sem}$ is a hyperparameter, and $\hat{S}$ and $S$ are the predicted and ground truth segmentation labels.
5 Experiments
Implementation details. We implemented FeatSplat with both $16-$ and $32-$dimensional feature vectors for novel view synthesis, and only $32-$dimensional ones for novel view synthesis plus semantic segmentation. For each scene, we train a MLP with one hidden layer of 64 neurons, common for all Gaussians. For semantic segmentation, the model is trained on 64 classes: the 63 most frequent classes in ScanNet++, plus an unknown category label. Further implementation and optimization details are included in App. A.
Baselines. We compare FeatSplat against 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis], the current state of the art for real-time novel view synthesis, Compact-3DGS [Lee et al.(2023)Lee, Rho, Sun, Ko, and Park] which also augmented 3DGS with learned feature vectors instead of SHs, and iNGP [Müller et al.(2022)Müller, Evans, Schied, and Keller], which was the previous state of the art. Finally, we also include Zip–NeRF and Mip-NeRF 360 metrics on the datasets they have been reported, as a reference. Nevertheless, FeatSplat should not be compared against them due to their offline nature (they both render at less than 1 frame per second).
Datasets and metrics. Rendering performance is evaluated using PSNR, LPIPS [Zhang et al.(2018)Zhang, Isola, Efros, Shechtman, and Wang], and SSIM [Wang et al.(2004)Wang, Bovik, Sheikh, and Simoncelli], computational aspects by the representation size in megabytes (MB), rendering speed in Frames per Second (FPS) on a GTX 3090, and semantic segmentation by the mIOU weighted over classes’ support among all scenes. We evaluate on three common datasets in the literature [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis, Lee et al.(2023)Lee, Rho, Sun, Ko, and Park, Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman]: Mip360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] (T&Ts); and Deep Blending [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow] (DB), for a total of 12 scenes. In these datasets the test set is selected by sampling one out of every eight frames of the image sequences, and the other 7 frames are used to train the model. We also evaluate FeatSplat in a subset of 21 scenes from the novel ScanNet++ [Yeshwanth et al.(2023)Yeshwanth, Liu, Nießner, and Dai]. ScanNet++ is a large–scale dataset with high-resolution RGB and semantic annotations of indoor scenes, and a test set generated from an independent camera trajectory specifically designed to evaluate novel view synthesis with low overlap and separated views.
Figure 2: Comparison of 3D position of test (red) and training (blue) points of view on one scene for each of the four datasets.
The independent test set from ScanNet++ represents a considerable increase in difficulty with respect to commonly used datasets. Sampling one out of 8 frames from a video stream (as done in Mip-360, T&Ts and DB) means that at least the previous and following frames would have a significant overlap with the test frames. In contrast, an independent sequence from a different trajectory results in test images looking at unexplored parts of the scene or at the same areas but from rather different viewpoints. In Figure 2 we can visualize, for each of the four datasets, the scene with the highest average angular distance from each test viewpoint (red) to its 10 closest training viewpoints (blue). The contrast between datasets is clear. While on Mip-360, T&Ts, and DB, cameras tend to be nearby each other and looking in the same direction, on ScanNet++ there are several test views that are isolated and looking at mostly unexplored areas. Further details on the datasets can be found in App. B.
Figure 3: Comparison of our method trained with features with 16 (FeatSplat–16) and 32 dimensions (FeatSplat–32), against prior work. From top to down the scenes are Bicycle, Bonsai, and Room from Mip-360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], and Playroom from DB [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow].
Figure 4: Novel view synthesis with low overlap of FeatSplat–32 against 3DGS. From left to right the scenes are Train from T&Ts [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun], Treehill, and Bicycle from Mip-360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman].
Mip-360 Tanks and Temples Deep Blending
SSIM PSNR LPIPS FPS Mem SSIM PSNR LPIPS FPS Mem SSIM PSNR LPIPS FPS Mem
INGP–Big 0.699 25.59 0.331 9.4 48MB 0.745 21.92 0.305 14.4 48MB 0.817 24.96 0.390 2.8 48MB
3D Gaussian Splatting 0.813 27.39 0.218 115 880MB 0.844 23.65 0.179 156 431MB 0.899 29.46 0.247 126 662MB
Compact-3DGS 0798 26.97 0.244 140 49MB 0.831 23.32 0.201 185 39MB 0.901 29.79 0.258 181 43MB
FeatSplat–16 (Ours) 0.808 27.48 0.230 96 275MB 0.848 24.48 0.185 171 149MB 0.904 29.78 0.249 111 212MB
FeatSplat–32 (Ours) 0.808 27.53 0.227 82 420MB 0.848 24.53 0.183 147 227MB 0.904 29.85 0.247 98 327MB
Mip–NeRF 360 0.792 27.29 0.237 0.06 9MB 0.759 22.22 0.257 0.14 9MB 0.901 29.4 0.245 0.09 9MB
ZipNeRF 0.836 28.54 0.177 0.25
Table 1: Evaluation on Mip-360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], T&Ts [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] and DB [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow] for novel view synthesis.
5.1 Novel View Synthesis Results
Mip-360, T&Ts and DB. The quantitative results on Mip-360 dataset, T&Ts and DB in Table 1 show how both FeatSplat models achieve similar rendering performance. Although our 32-dimensions variant has a slight advantage on rendering quality (equal or better PSNR, LPIPS and SSIM), the 16-dimensional version has higher rendering speed and a smaller representation size. We conclude that 16 dimensions is enough to encode a scene’s RGB information. Comparing FeatSplat against 3DGS and Compact-3DGS, our method consistently achieves the best PSNR on all three datasets, with a better SSIM on two of them. Nevertheless, 3DGS has a better LPIPS overall. The qualitative results in Figure 3 show how all methods render high quality images, although with slight differences. 3DGS generates more accurate high-frequency details due to its tendency to over-densify scenes, e.g. the grass on the Bicycle scene and the trees on the Train scene. Instead, FeatSplat–32 is also able to capture details missed by 3DGS (e.g. the bush in Bicycle, the door in Bonsai, and the floor in Playroom) while reducing the amount of artifacts thanks to its more compact and less overfitted representation. This compactness can also be observed in the memory used, that is half for FeatSplat–32 and a fourth for FeatSplat–16. It can be noticed that the rendering speed is reduced for our FeatSplat in comparison to 3DGS and Compact-3DGS. However, we are still well within real-time limits, so we believe this is not a significant limitation.
ScanNet++
SSIM PSNR LPIPS FPS Mem
3D Gaussian Splatting 0.867 23.83 0.250 177 217
Compact 3DGS 0.854 22.76 0.272 170 35
FeatSplat–16 (Ours) 0.875 24.62 0.245 82 74
FeatSplat–32 (Ours) 0.875 24.62 0.245 71 112
Table 2: Features Splatting evaluation on 21 scenes from ScanNet++
Out of distribution. In order to analyze points of view far from the training cameras, we sample four points of view from simple linear trajectories on outdoor scenes, see Figure 4. Observe how in Train 3DGS generates floating 3D Gaussians to model far away clouds, which obfuscate the visibility for closer views. Observe instead in FeatSplat–32 that, as the camera gets closer to the train, the Gaussians that before were white, first switch to blue and finally also to green, to better represent the background sky and trees. A similar effect can be observed on the Bicycle scene in the right top corner of the image. Finally, the Treehill data reveals how FeatSplat learned to adapt the Gaussians’ brightness for slightly different positions of the camera while 3DGS did not.
Figure 5: Novel view synthesis of FeatSplat–16 and FeatSplat–32, against 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] and Compact-3DGS [Lee et al.(2023)Lee, Rho, Sun, Ko, and Park] on a subset of scenes from ScanNet++ [Yeshwanth et al.(2023)Yeshwanth, Liu, Nießner, and Dai].
ScanNet++. Table 2 shows that FeatSplat significantly outperforms both 3DGS and Compact-3DGS on all SSIM, PSNR, and LPIPS, with both FeatSplat–16 and FeatSplat–32 achieving the same performance. The gap between FeatSplat and 3DGS is due to a more challenging test set, as previously mentioned. To properly learn view-dependent effects, 3DGS overfits to the training cameras. Some Gaussians are placed in such positions that are seen only in the optimized direction, while not being visible from other training directions without the same effect. This effect is unnoticed from points of view that are close to the training samples, or with a high visual overlap. However, for viewpoints far-away from the training ones, or with lower visual overlap, they become easily noticeable. The ability of FeatSplat to encode multiple colors in the same Gaussian, and to condition the decoding on the point of view, reduces the impact of this problem. Figure 5 shows how FeatSplat renders more accurate colors, captures smaller details, and reduces the number of artifacts.
Figure 6: Semantic FeatSplat with 32-dimensional vectors and 64 classes on ScanNet++.
Semantic FeatSplat Our results so far suggest that 16 dimensions suffice to encode color. We then chose FeatSplat–32 to encode both RGB and semantic labels and perform semantic segmentation. FeatSplat’s segmentation masks for novel views achieve weighted mIOU of $0.629$, maintaining the rendering performance with a PSNR of $24.64$, and SSIM of $0.875$ and an LPIPS of $0.244$, and still rendering in real-time at 56 FPS. The segmentation maps generated, see Figure 6, show that the model produces coherent segmentations even for small elements. However, the edges of the masks are noisy, as they follow Gaussians’ shapes.
5.2 Limitations
The small size of our MLP allows real-time rendering but limits the complexity of textures encoded in each Gaussian. Therefore, our FeatSplat still generates small artifacts in unobserved areas. Furthermore, our model sometimes over-smooths some highly textured areas, like the grass or some tree leaves. This trade-off between capacity and speed can have a negative impact in applications focused on rendering around the training cameras, for which generalization is less relevant. Finally, FeatSplat is slower than 3DGS in part due to the CUDA kernel implementation designed for RGB channels, with rendering time scaling linearly with the dimension of the alpha-blended feature vector. Nevertheless, the increased capacity of FeatSplat vectors halves the number of Gaussians per scene, limiting the training and rendering slow-down to just a half of 3DGS at a resolution of $1752\times 1168$.
6 Conclusions
In this paper we showed how the use of feature vectors to encode color information surpasses the performance of spherical harmonics when linked to small 3D primitives such as Gaussians for novel view synthesis. We argue that SHs lack capacity, and as a consequence an overpopulation of 3D Gaussians is required in order to approximate complex textures, which in turns results in the model overfitting the training viewpoints. Our FeatSplat overcomes this issue by using feature vectors, alpha blended, concatenated to camera parameters and decoded by an MLP, that render a higher variety of more complex textures and hence reduce the Gaussians’ overpopulation. FeatSplat outperforms previous methods on standard datasets in the literature and on the challenging ScanNet++ [Yeshwanth et al.(2023)Yeshwanth, Liu, Nießner, and Dai]. In addition, we keep a competitive real-time rendering and enable the extension to other tasks. Finally, it is worth mentioning that FeatSplat should also benefit from any 3DGS optimizations, like the quantization proposed by Compact-3DGS [Lee et al.(2023)Lee, Rho, Sun, Ko, and Park], which would speed up its rendering.
Appendix A Model details
A.1 Implementation
We use a two layer MLP with bias. The first layer has 64 neurons and a SiLU activation function [Hendrycks and Gimpel(2023)]. For the only RGB implementation, the output layers have 3 output neurons and a Sigmoid activation function. Instead, for the semantic segmentation implementation, the output layer has 67 output neurons, with a sigmoid activation function for the first 3 channels, and a Softmax for the last 64. We use the Adam optimizer with $\epsilon=1\times 10^{-15}$, a learning rate of $0.001$ for the MLP and $0.0025$ for the feature vectors, and $\lambda_{sem}=0.001$ for the semantic segmentation model. The rest of the Gaussians’ parameters and optimization hyperparameters follow 3DGS’s [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] original implementation. We use the 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] CUDA differentiable rasterizer. We adapt it to our model by skipping spherical harmonics and passing feature vectors as precomputed colors. To do it, we change the number of channels when compiling from 3 to 16 and 32 for FeatSplat-16 and FeatSplat-32 respectively. The model has been implemented on Pytorch 1.12 and CUDA SDK 11.6.2.
A.2 Embeddings ablation
We ablate the impact on the 32-dimensional model of concatenating to the pre-MLP feature vectors the 3D camera postion coordinates $\mathbf{x}_{cam}$, its rotation using Euler angles $\mathbf{x}_{rot}$, and the 2 dimension embedding of the relative position of each pixel $\mathbf{e}_{p}$. The evaluation is done on the scenes Bicycle from Mip-360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman] and Train from T&T  [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun].
Bicycle Train
FeatSplat – 32 SSIM PSNR LPIPS SSIM PSNR LPIPS
w/o embeddings 0.7415 24.54 0.241 0.807 22.02 0.221
only $\mathbf{e}_{p}$ 0.745 24.70 0.239 0.815 22.84 0.216
only $\mathbf{x}_{cam}$ 0.748 24.84 0.239 0.808 22.09 0.219
only $\mathbf{x}_{rot}$ 0.743 24.52 0.240 0.810 22.03 0.22
with $\mathbf{e}_{p}$ and $\mathbf{x}_{rot}$ 0.746 24.69 0.239 0.813 22.748 0.215
with $\mathbf{x}_{cam}$ and $\mathbf{x}_{rot}$ 0.748 24.75 0.240 0.813 22.58 0.216
with $\mathbf{e}_{p}$ and $\mathbf{x}_{cam}$ 0.751 25.02 0.237 0.815 22.85 0.215
with $\mathbf{e}_{p}$, $\mathbf{x}_{cam}$ and $\mathbf{x}_{rot}$ 0.749 25.01 0.238 0.813 22.84 0.216
Table 3: FeatSplat embbedings ablation on scenes Bicycle from Mip-360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman] and Train from T&T  [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun]. We integrate in the final model only $\mathbf{e}_{p}$ and $\mathbf{x}_{cam}$.
The ablation results in Table 3 show that both $\mathbf{e}_{p}$ and $\mathbf{x}_{cam}$ have by themselves a positive impact on the metrics, while $\mathbf{x}_{cam}$ does not improve against not adding it. While $\mathbf{e}_{p}$ is responsible from almost all the improvement on Train, $\mathbf{x}_{cam}$ has a better performance on Bicycle, and both together achieve the best performance. On the other hand, $\mathbf{x}_{rot}$ alone does not improve the performance. Combining it with $\mathbf{x}_{cam}$ improves the performance of $\mathbf{x}_{cam}$ alone on Train, although less than $\mathbf{e}_{p}$. The results show that $\mathbf{x}_{cam}$ is redundant, with $\mathbf{e}_{p}$ and $\mathbf{x}_{cam}$ having the best performance.
Appendix B Datasets
B.1 Mip360, T&T and DB.
The three typical datasets we use amount to a total of 12 scenes: Mip360[Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], with 9 scenes, 5 outdoor at $1237\times 822$ and 4 indoor at $1558\times 1039$, with an average of 216 views per scene; Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] with 2 outdoor scenes at $979\times 546$ and an average of 276 views per scene; and Deep Blending [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow] with 2 indoor scenes at $1264\times 832$ and an average of 244 views per scene. In these datasets the testing set is selected sampling one every eight frames of the camera trajectory, using the other 7 frames to train.
B.2 ScanNet++
ScanNet++ is large-scale dataset with 460 indoor scenes and high-quality RGB and semantic annotations. It has been designed for evaluating high resolution novel view synthesis, with a resolution of undistorted images of $1752\times 1168$ and a test set generated from an independent camera trajectory. The subset of used scenes has been randomly selected ensuring the fit in memory 24 gigabytes of vRAM in a single RTX 3090 GPU as Float-Point 32. As a consequence, the selected subset of scenes has an average of 408 views per scene, higher than the others datasets used. The evaluated scenes are: 0a5c013435, f07340dfea, 7bc286c1b6, d2f44bf242, 85251de7d1, 0e75f3c4d9, 98fe276aa8, 7e7cd69a59, f3685d06a9, 21d970d8de, 8b5caf3398, ada5304e41, 4c5c60fa76, ebc200e928, a5114ca13d, 5942004064, 1ada7a0617, f6659a3107, 1a130d092a, 80ffca8a48, 08bbbdcc3d,
B.3 Test and Train cameras plots
Figures 879 and 10 show the distribution of train and test cameras for all evaluated scenes.
Figure 7: Train (blue) and test (red) and camera for ScanNet++
Figure 8: Train (blue) and test (red) and camera for ScanNet++
Figure 9: Train (blue) and test (red) and camera for T&T and DB
Figure 10: Train (blue) and test (red) and camera for Mip-360

### v2

@@
 Finally, we also show the capacity and convenience of our feature vector representation, demonstrating its capability not only to generate RGB values for novel views, but also their per-pixel semantic labels.
-We will release the code upon acceptance.
+Code is available on the project page.
 Keywords: Gaussian Splatting, Novel View Synthesis, Feature Splatting
@@
 In general, representations should not only encode a scene’s geometry.
-Radiance properties are very relevant for novel view synthesis and high-level aspects such as semantics or affordances are essential to define robotics tasks.
+Radiance properties are very relevant for novel view synthesis, and high-level aspects such as semantics or affordances are essential to define robotics tasks.
 In the last years, Neural Radiance Fields (NeRFs) [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] have been the dominant approach in the research literature to learn implicit scene representations from images.
@@
 LangSplat [Qin et al.(2023)Qin, Li, Zhou, Wang, and Pfister] combines CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] features with SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] segmentation masks to generate a language field enabling open-vocabulary 2D semantic segmentation.
+Similarly, both Feature 3DGS [Zhou et al.(2024)Zhou, Chang, Jiang, Fan, Zhu, Xu, Chari, You, Wang, and Kadambi] and Gaussian Grouping [Ye et al.(2024)Ye, Danelljan, Yu, and Ke], use SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] as a teacher to embed in each Gaussian a semantic feature.
 FMGS [Zuo et al.(2024)Zuo, Samangouei, Zhou, Di, and Li] uses a hash grid based on iNGP [Müller et al.(2022)Müller, Evans, Schied, and Keller] to encode CLIP and DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, Howes, Huang, Xu, Sharma, Li, Galuba, Rabbat, Assran, Ballas, Synnaeve, Misra, Jegou, Mairal, Labatut, Joulin, and Bojanowski] features, to also generate an intrinsic representation capable of open-vocabulary 2D segmentation.
@@
 Differentiable rendering.
-To render an RGB image $\hat{{I}}\in\mathbb{R}^{w\times h\times 3}$ of the scene, first SHs are converted to RGB, $\mathbf{c}_{G}=\mathcal{S}_{\mathbf{k}_{G}}\left(\mathbf{d}_{G}\right)$.
+To render a RGB image $\hat{{I}}\in\mathbb{R}^{w\times h\times 3}$ of the scene, first SHs are converted to RGB, $\mathbf{c}_{G}=\mathcal{S}_{\mathbf{k}_{G}}\left(\mathbf{d}_{G}\right)$.
 Then, Gaussians are projected, or splatted, into 2D Gaussians in the image plane [Zwicker et al.(2001)Zwicker, Pfister, Van Baar, and Gross] and sorted using the Radix algorithm.
@@
 where $\lambda_{sem}$ is a hyperparameter, and $\hat{S}$ and $S$ are the predicted and ground truth segmentation labels.
+Figure 2: Comparison of 3D position of test (red) and training (blue) points of view on one scene for each of the four datasets.
 5 Experiments
@@
 Nevertheless, FeatSplat should not be compared against them due to their offline nature (they both render at less than 1 frame per second).
+Finally, on the semantic segmentation experiment, we do not compare against other models that learn specific semantic features for open-vocabulary semantic segmentation.
+Given that our model is trained on a closed-vocabulary, we would have an unfair advantage.
+Figure 3: Comparison of our method trained with features with 16 (FeatSplat–16) and 32 dimensions (FeatSplat–32), against prior work.
+From top to down the scenes are Bicycle, Bonsai, and Room from Mip-360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], and Playroom from DB [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow].
 Datasets and metrics.
@@
 ScanNet++ is a large–scale dataset with high-resolution RGB and semantic annotations of indoor scenes, and a test set generated from an independent camera trajectory specifically designed to evaluate novel view synthesis with low overlap and separated views.
-Figure 2: Comparison of 3D position of test (red) and training (blue) points of view on one scene for each of the four datasets.
 The independent test set from ScanNet++ represents a considerable increase in difficulty with respect to commonly used datasets.
@@
 Further details on the datasets can be found in App.
-B.
-Figure 3: Comparison of our method trained with features with 16 (FeatSplat–16) and 32 dimensions (FeatSplat–32), against prior work.
-From top to down the scenes are Bicycle, Bonsai, and Room from Mip-360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], and Playroom from DB [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow].
+C.
 Figure 4: Novel view synthesis with low overlap of FeatSplat–32 against 3DGS.
@@
 3D Gaussian Splatting 0.813 27.39 0.218 115 880MB 0.844 23.65 0.179 156 431MB 0.899 29.46 0.247 126 662MB
-Compact-3DGS 0798 26.97 0.244 140 49MB 0.831 23.32 0.201 185 39MB 0.901 29.79 0.258 181 43MB
+Compact-3DGS 0.798 26.97 0.244 140 49MB 0.831 23.32 0.201 185 39MB 0.901 29.79 0.258 181 43MB
 FeatSplat–16 (Ours) 0.808 27.48 0.230 96 275MB 0.848 24.48 0.185 171 149MB 0.904 29.78 0.249 111 212MB
@@
 5.1 Novel View Synthesis Results
+Figure 5: Global illumination change modifying MLP input values of the pixel embedding (PE), and the 3D coordinates (X, Y, Z), on scene 03f7a0e617 from ScanNet++.
 Mip-360, T&Ts and DB.
@@
 In order to analyze points of view far from the training cameras, we sample four points of view from simple linear trajectories on outdoor scenes, see Figure 4.
-Observe how in Train 3DGS generates floating 3D Gaussians to model far away clouds, which obfuscate the visibility for closer views.
+Observe how, in Train, 3DGS generates floating 3D Gaussians to model far away clouds, which obfuscate the visibility for closer views.
 Observe instead in FeatSplat–32 that, as the camera gets closer to the train, the Gaussians that before were white, first switch to blue and finally also to green, to better represent the background sky and trees.
@@
 Finally, the Treehill data reveals how FeatSplat learned to adapt the Gaussians’ brightness for slightly different positions of the camera while 3DGS did not.
+Light encoding.
+Furthermore, we analyze the pixel and position embeddings impact on the MLP by individually evaluating different input values for each of them, see Figure 5.
+Apparently, the MLP learns to use these embeddings to codify global lightning information.
+As a consequence, our model allows to modify the scene lightning at inference time without any addition training.
 Figure 6: Novel view synthesis of FeatSplat–16 and FeatSplat–32, against 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] and Compact-3DGS [Lee et al.(2023)Lee, Rho, Sun, Ko, and Park] on a subset of scenes from ScanNet++ [Yeshwanth et al.(2023)Yeshwanth, Liu, Nießner, and Dai].
@@
 Figure 6 shows how FeatSplat renders more accurate colors, captures smaller details, and reduces the number of artifacts.
-Figure 6: Semantic FeatSplat with 32-dimensional vectors and 64 classes on ScanNet++.
-Semantic FeatSplat Our results so far suggest that 16 dimensions suffice to encode color.
+Semantic FeatSplat
+Our results so far suggest that 16 dimensions suffice to encode color.
 We then chose FeatSplat–32 to encode both RGB and semantic labels and perform semantic segmentation.
@@
 However, the edges of the masks are noisy, as they follow Gaussians’ shapes.
+Nevertheless, it proves the versatility of these features and their potential to encode different kind of information.
+Figure 7: Semantic FeatSplat with 32-dimensional vectors and 64 classes on ScanNet++.
 5.2 Limitations
@@
 A.1 Implementation
-We use a two layer MLP with bias.
+We use a two-layer MLP with bias.
 The first layer has 64 neurons and a SiLU activation function [Hendrycks and Gimpel(2023)].
-For the only RGB implementation, the output layers have 3 output neurons and a Sigmoid activation function.
+For the RGB-only implementation, the output layers have 3 output neurons and a Sigmoid activation function.
 Instead, for the semantic segmentation implementation, the output layer has 67 output neurons, with a Sigmoid activation function for the first 3 channels, and a Softmax for the last 64.
@@
 Combining it with $\mathbf{x}_{cam}$ improves the performance of $\mathbf{x}_{cam}$ alone on Train, although less than $\mathbf{e}_{p}$.
-The results show that $\mathbf{x}_{cam}$ is redundant, with $\mathbf{e}_{p}$ and $\mathbf{x}_{cam}$ having the best performance.
-Appendix B Datasets
-B.1 Mip360, T&T and DB.
-The three typical datasets we use amount to a total of 12 scenes: Mip360[Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], with 9 scenes, 5 outdoor at $1237\times 822$ and 4 indoor at $1558\times 1039$, with an average of 216 views per scene; Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] with 2 outdoor scenes at $979\times 546$ and an average of 276 views per scene; and Deep Blending [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow] with 2 indoor scenes at $1264\times 832$ and an average of 244 views per scene.
+The results show that $\mathbf{x}_{rot}$ is redundant, with $\mathbf{e}_{p}$ and $\mathbf{x}_{cam}$ having the best performance.
+Appendix B ScanNet++ Novel View Synthesis Benchmark
+ScanNet++
+SSIM PSNR LPIPS FPS
+HW GTS - RPBG 0.873 24.355 0.280 $\thicksim 1$
+FeatSplat–32 (Ours) 0.869 24.247 0.314 50
+Nerfacto 0.861 24.049 0.342 0.3
+TensoRF 0.849 23.978 0.407
+3D Gaussian Splatting 0.871 23.891 0.319 155
+iNGP 0.859 23.812 0.375 0.3
+Table 4: ScanNet++ Novel View Synthesis Benchmark
+Appendix C Datasets
+C.1 Mip360, T&T and DB.
+The three typical datasets we use amount to a total of 12 scenes: Mip360[Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], with 9 scenes, 5 outdoor at $1237\times 822$ (Bicycle, Flowers, Garden, Stump, Treehill,) and 4 indoor at $1558\times 1039$ (Bonsai, Counter, Kitchen, Room), with an average of 216 views per scene; Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] with 2 outdoor scenes at $979\times 546$ (Train and Truck) and an average of 276 views per scene; and Deep Blending [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow] with 2 indoor scenes (DrJohnson and Playroom) at $1264\times 832$ and an average of 244 views per scene.
 In these datasets the testing set is selected sampling one every eight frames of the camera trajectory, using the other 7 frames to train.
-B.2 ScanNet++
+C.2 ScanNet++
 ScanNet++ is large-scale dataset with 460 indoor scenes and high-quality RGB and semantic annotations.
@@
 The evaluated scenes are: 0a5c013435, f07340dfea, 7bc286c1b6, d2f44bf242, 85251de7d1, 0e75f3c4d9, 98fe276aa8, 7e7cd69a59, f3685d06a9, 21d970d8de, 8b5caf3398, ada5304e41, 4c5c60fa76, ebc200e928, a5114ca13d, 5942004064, 1ada7a0617, f6659a3107, 1a130d092a, 80ffca8a48, 08bbbdcc3d,
-B.3 Test and Train cameras plots
-Figures 879 and 10 show the distribution of train and test cameras for all evaluated scenes.
+C.3 Test and Train cameras plots
+Figures 9810 and 11 show the distribution of train and test cameras for all evaluated scenes.
 Figure 8: Train (blue) and test (red) and camera for ScanNet++